OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. P(6, 2), Q(8, -4)

OpenStudy (anonymous):

so we need to find a line of the form y = mx+b m is the slope (m = (y2-y1)/(x2-x1)) b is the y intercept. the y value when x = 0. you are given two points P(x1,y1) = P(6,2) and Q(x2,y2) = Q(8,-4). solve for the slope first, then we can work through finding the y intercept.

OpenStudy (anonymous):

my answer has to be _X_Y and then what ever comes after Y

OpenStudy (anonymous):

well once we solve the y = mx + b form we can just switch around the numbers to get y -mx = b.

OpenStudy (anonymous):

I dont got a M or b to put in there like i said it has to be __ X__Y and then what ever comes after y

OpenStudy (anonymous):

show me an example

OpenStudy (anonymous):

I don't really have one to show u

OpenStudy (anonymous):

well is it the point slope form or slope intercept form.

OpenStudy (anonymous):

all it give me is "Indicate in standard form the equation of the line passing through the given points." P(6, 2), Q(8, -4)